1. what is the y-intercept of the linear equation y= -3x?
2. what is the slope of the equation y= -4?
3. rearrange x + y= 4 into slope-intercept,and identify the slope and y-intercept?
4. given a slope of 5/3 and a y-intercept of 3, write the equation of the line in y= mx+b?
5. Name one point(other than the y-intercept) that is on the line with the given equation?

Answer by stanbon(75887)

1. what is the y-intercept of the linear equation y= -3x?
Ans: (0,0)
------------------
2. what is the slope of the equation y= -4?
Ans: m = -4
------------------
3. rearrange x + y= 4 into slope-intercept,and identify the slope and y-intercept?
Ans: y = -x+4
slope = -1
y-intercept: 4
--------------------------
4. given a slope of 5/3 and a y-intercept of 3, write the equation of the line in y= mx+b?
Ans: y = (5/3)x + 3
--------------------------
5. Name one point(other than the y-intercept) that is on the line with the given equation?
Ans: If x = 3, y = 5 + 3 = 8
Point: (3,8)
===================
Cheers,
Stan H.
===================